Parth (parthkohli):

Just solved this problem. Give it a try: How many integers \(1≤N≤1000\) can be written both as the sum of \(26\) consecutive integers and as the sum of \(13\) consecutive integers?

Parth (parthkohli):

It's very easy!

OpenStudy (experimentx):

use these conditions 13(2a+25) < 1000 6(2a+12) < 1000 13(2a+25) = 6(2a+12)

Parth (parthkohli):

So what do you get as your answer?

OpenStudy (experimentx):

just plugin to W|A

OpenStudy (experimentx):

seems the result is negative

Parth (parthkohli):

@FoolAroundMath got the correct answer (in the messages).

Parth (parthkohli):

So I can disclose the solution.

Parth (parthkohli):

FoolAroundMath is the answer to your question 38? Yup!

OpenStudy (experimentx):

ah .. aren't these you 13 consecutive numbers 84 96 108 120 132 144 156 168 180 192 204 216 228 240 Columns 15 through 28 252 264 276 288 300 312 324 336 348 360 372 384 396 408 Columns 29 through 42 420 432 444 456 468 480 492 504 516 528 540 552 564 576 Columns 43 through 56 588 600 612 624 636 648 660 672 684 696 708 720 732 744 Columns 57 through 70 756 768 780 792 804 816 828 840 852 864 876 888 900 912 Columns 71 through 84 924 936 948 960 972 984 996 84 96 108 120 132 144 156 Columns 85 through 98 168 180 192 204 216 228 240 252 264 276 288 300 312 324 Columns 99 through 112 336 348 360 372 384 396 408 420 432 444 456 468 480 492 Columns 113 through 126 504 516 528 540 552 564 576 588 600 612 624 636 648 660 Columns 127 through 140 672 684 696 708 720 732 744 756 768 780 792 804 816 828 Columns 141 through 154 840 852 864 876 888 900 912 924 936 948 960 972 984 996

Parth (parthkohli):

Huh

OpenStudy (experimentx):

looks like bug in my program.

OpenStudy (shubhamsrg):

@ParthKohli n+(n+1) ..(n+26) ---> these are 27 terms and not 26 ..!!

Parth (parthkohli):

Actually \(25\)... haha sorry

Parth (parthkohli):

But my answer is still correct. That's for sure :)

OpenStudy (shubhamsrg):

yes i get it,its still correct..

OpenStudy (shubhamsrg):

and not 25,they are, 27 terms only.. o.O

Parth (parthkohli):

\[N = n + (n + 1) + (n + 2)\cdots (n + 25) = 26(n + \text{something)}\]and\[N = k + ( k+1) + (k + 2)\cdots (k+12)=13(k + \text{something)}\]So, the number must be a multiple of 13 and 26. If a number is a multiple of 26, it is a multiple of 13. So we have to find the number of multiples of 26 less than or equal to 1000 which is \(\left\lfloor1000/26\right\rfloor=38\).

Parth (parthkohli):

@shubhamsrg I said that it'd be \(n + (n + 1)\cdots(n + 25)\)

OpenStudy (shubhamsrg):

ohh.. :P

Parth (parthkohli):

:)

OpenStudy (experimentx):

here's a short code for doing that on mathematica Length[Intersection[Table[Sum[j, {j, i, i + 12}], {i, 1, 70}], Table[Sum[j, {j, i, i + 25}], {i, 1, 22}]]]